# Bipartite Diophantine tuples and their applications

### v1

@@ 1. Introduction
It is shown in [32, Theorem 2.3] that, for Question 1.3, we have $\ell\leq 9$ for $k=3$, $\ell\leq 6$ for $k=4$, $\ell\leq 5$ for $k=5$, and $\ell\leq 4$ for $k\geq 6$. However, as we discussed earlier, when $k=2$, we do not know any upper bound on $\ell$. In [32, Remark 4.9], the second author asked if some improved bounds can be proved in Question 1.3 under the ABC conjecture. We answer this question in the affirmative in Theorem 1.4 below; in particular, we show that $\ell=2$ if $k\geq 6$, $\ell\leq 3$ if $k\in\{4,5\}$, and $\ell\leq 5$ if $k=3$.
@@
We remark that there are some known bounds for Question 1.3 under other conjectures. As remarked in [32, Remark 4.9], if $k\geq 5$, we can predict from the uniformity conjecture [10] that the answer to Question 1.3 is $\ell=2$ by considering superelliptic curves of the form $C:y^{k}=(a_{1}x+n)(a_{2}x+n)$ with $a_{1}\neq a_{2}$. Similarly, using the uniformity conjecture, for $k\in\{3,4\}$, we can predict that $\ell\leq 3$ by considering superelliptic curves of the form $C:y^{k}=(a_{1}x+n)(a_{2}x+n)(a_{3}x+n)$ with $a_{1},a_{2},a_{3}$ distinct; for $k=2$, we can predict that $\ell\leq 5$ by considering hyperelliptic curves of the form
@@
Our proofs of these new results combine various tools from additive combinatorics, Diophantine approximation, Diophantine equations, and sieve methods. Some of the ideas are inspired by [5, 11, 16, 32].
@@ 2.1. A gap principle and its applications
The following proposition is a slightly weaker version of [32, Proposition 4.1]. Its proof is based on an explicit bound on the number of solutions of Thue inequalities by Evertse [19] combined with repeated applications of Lemma 2.1.
Proposition 2.4 ([32, Proposition 4.1]).
@@ 2.2. Applications of sieve methods
Then, when $N$ is sufficiently large compared to $k,m,C,L$ so that $Q$ is sufficiently large, estimates (1) and (2) imply that
@@ Acknowledgments
The authors thank Michael Bennett, Ernie Croot, Andrej Dujella, Greg Knapp, and Attila Pethő for helpful discussions. The first author was supported in part by an NSERC Discovery Grant. The second author was supported in part by an NSERC fellowship.

### v2

@@ 1. Introduction
It is shown in [31, Theorem 2.3] that, for Question 1.3, we have $\ell\leq 9$ for $k=3$, $\ell\leq 6$ for $k=4$, $\ell\leq 5$ for $k=5$, and $\ell\leq 4$ for $k\geq 6$. However, as we discussed earlier, when $k=2$, we do not know any upper bound on $\ell$. In [31, Remark 4.9], the second author asked if some improved bounds can be proved in Question 1.3 under the ABC conjecture. We answer this question in the affirmative in Theorem 1.4 below; in particular, we show that $\ell=2$ if $k\geq 6$, $\ell\leq 3$ if $k\in\{4,5\}$, and $\ell\leq 5$ if $k=3$.
@@
We remark that there are some known bounds for Question 1.3 under other conjectures. As remarked in [31, Remark 4.9], if $k\geq 5$, we can predict from the uniformity conjecture [10] that the answer to Question 1.3 is $\ell=2$ by considering superelliptic curves of the form $C:y^{k}=(a_{1}x+n)(a_{2}x+n)$ with $a_{1}\neq a_{2}$. Similarly, using the uniformity conjecture, for $k\in\{3,4\}$, we can predict that $\ell\leq 3$ by considering superelliptic curves of the form $C:y^{k}=(a_{1}x+n)(a_{2}x+n)(a_{3}x+n)$ with $a_{1},a_{2},a_{3}$ distinct; for $k=2$, we can predict that $\ell\leq 5$ by considering hyperelliptic curves of the form
@@
Question 1.6 essentially asks how to measure the largest possible multiplicative complexity of the set of shifted $k$-th powers. We expect that the maximum possible dimension of multiplicative Hilbert cubes is much smaller than that of bipartite Diophantine tuples or $P_{n}^{(k)}$ sets, as the defining conditions for multiplicative Hilbert cubes impose substantially more multiplicative constraints.
@@
Our proofs of these new results combine various tools from additive combinatorics, Diophantine approximation, Diophantine equations, and sieve methods. Some of the ideas are inspired by [5, 11, 16, 31]. Our proofs of the above results related to Hilbert cubes crucially depend on the many multiplicative constraints imposed by the defining conditions of multiplicative Hilbert cubes.
@@ 2.1. A gap principle and its applications
The following proposition is a slightly weaker version of [31, Proposition 4.1]. Its proof is based on an explicit bound on the number of solutions of Thue inequalities by Evertse [19] combined with repeated applications of Lemma 2.1.
Proposition 2.4 ([31, Proposition 4.1]).
@@ 2.2. Applications of sieve methods
Then, when $N$ is sufficiently large compared to $k,m,C,L,\delta$ so that $Q$ is sufficiently large, estimates (1) and (2) imply that
@@ Acknowledgments
The authors thank Michael Bennett, Ernie Croot, Andrej Dujella, Greg Knapp, and Attila Pethő for helpful discussions. The authors also thank the anonymous referees for their valuable comments and suggestions. The first author was supported in part by an NSERC Discovery Grant. The second author was supported in part by an NSERC fellowship.